the areA of square plate is calculated by measuring the length to be 5.5 cm. if the scale used has the leAST COUNT OF 0.1 CM THEN WHAT IS THE MAXIMUM PERCENTAGE ERROR IN THE AREA?
 

Grade:11

1 Answers

Dhruv
6 years ago
Area = length X length    (because of square)
A = L2  
∂A/A  = 2 (∂L/L)
% error of A = 2 (% error of A)
% error of A = 2 (0.1/5.5) X 100
% error of A = 200/55 %
% error of A = 40/11 % = 3.64 %